Security reported Saturday night that the connectivity to Trimble Hall had failed. Several of the residents also reported that the Internet connection had been lost.
Trouble shooting revealed that one of the two fiber strands between McIntyre and Trimble had failed. The service was moved to another pair of fibers and service was restored about 9:30 PM Saturday, April 26th.
SCYLD was promoted to domain controller in ACADEMIA. The procedure went very smoothly. About one hour after promotion, the RID attributes on the primary controller were updated, and the problem was cleared.
See Account Management in ACADEMIA Fails IV – Problem fixed (April 16, 2003) for a description of why this was done.
To gain space in Mc007, the servers listed above were moved to a new 4-post rack. This rack replaces the one that used to house the mail server and the DNS servers.
The latest Microsoft software updates were applied to MERLIN2 and VERONICA during PM.
In preparation for the promotion of SCYLD to domain controller, the latest Microsoft software patches were applied to it. These included Windows 2000 Service Pack 3, SQL Server 2000 Service Pack 3, and the latest IIS Security Rollup package, among others.
We attempted to reset RID attributes in Active Direcory for the ACADEMIA domain with the Java-based LDAP browser. Unfortunately, these attributes are read-only and can only be modified by the SYSTEM account.
We decided that we should bring another domain controller online in order to transfer or sieze RID master operations. It was my hope that the presence of another domain controller would force the RID master to issue another pool of RIDs, and clear the problem.
The only real candidate for promotion to domain controller was SCYLD. I obtained permission from Math/CS to promote SCYLD.
Josh made me aware of the problem. Jeremy found that the problem is due to the RID Operations Master reaching the end of its allocation pool. We tried deleting over 100 accounts. This did not fix the problem.
The Google search software was updated to version 3.4.12.
The Google system software was updated to version 1.16. Both updates address security issues.
During a modification of the syslog setting on the PIX firewall it hung, blocking all outbound/inbound network traffic for approximately 15 minutes. The rule base was determined to be correct, and the PIX was reset with a hard-reboot. Network traffic was restored.
Account creations in the ACADEMIA domain began failing. Attempts to create accounts or add computers to the domain hung the operations master for the domain. This caused slow roaming profile loading across the domain. It became necessary to reboot the operation masters.
Anti-SPAM rules on mail server modified to allow VPN clients to relay mail through the University mail server.
The main and mcintyre subnet sections of the dhcpd.conf file on the DHCP server were reconfigured to use porthos and alexandria for WINS instead of merlin. This is a change to test the new WINS configuration.
The /home volume on the mail server filled from 97% to 99% in about 6 hours. In order to prevent problems with the mail server, data was moved from the /home volume to /home3. The LDAP directory was updated and permissions were reassigned to the relocated data. The /home volume is currently at 75% capacity.
Automatic monthly Apache log rotation was set up on the main web server, www.ups.edu.